Milar Corporation makes a product with the following standard costs:

Standard Quantity or Hours   Standard Price or Rate

Direct materials 12.0 pounds $11.50 per pound

Direct labor 0.8 hours $36.00 per pound

Variable overhead 0.8 hours $17.00 per pound

In January the company produced 3,470 units using 13,880 pounds of the direct material and 2,896 direct labor-hours. During the month, the company purchased 16,900 pounds of the direct material at a cost of $14,640. The actual direct labor cost was $103,840 and the actual variable overhead cost was $47,220.

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The labor rate variance for January is:

A) 416F

B) 416U

C) 3,904U

D)3,904F
Answer #1

Solution:

Actual rate of labor = Actual direct labor cost / Actual labor hours = $103840 / 2896 = $35.856354

labor rate variance = (SR - AR) * Actual Hours = ($36 - $35.856354) * 2896 = $416 Favorable

Hence option "A" is correct.